# Eigenvalues of Heckman-Polychronakos operators

### v1

@@ 5.4. $a\dots ab\dots b$ case with $p=2$
Another situation where the general formula of Theorem 5.4 is fully explicit is $p=2$. In fact, in this special case we are able to compute all eigenvalues $\mathrm{eig}_{m}(\lambda,\tau,i)$ rather than only their sums over $i$. We fix $N=1,2,\dots$ and $0\leq\eta\leq N$ and set the multiplicities to be $(n_{1},n_{2})=(N-\eta,\eta)$. We further choose $a>b\geq 0$ and set
@@
We need some facts from the representation theory of symmetric groups, which can be found e.g. in [Dun78], or [Sta84, Section 3], or [CSST08, Sections 6.1, 6.2]. In particular, $V_{\lambda}$ expands into a direct multiplicity one sum of irreducible representations $\tau=(N-k,k)$, $0\leq k\leq\min(\eta,N-\eta)$. The operator $(T_{1}^{m}+\dots+T_{N}^{M})$ commutes with the action of symmetric group. Therefore, by the Schur’s lemma, it necessary acts as a multiple of identity in each $V_{\lambda;\tau}$ (which would not have been true if $\tau$ had multiplicity larger than $1$ in $\lambda$). Hence, the eigenvalues $\mathrm{eig}_{m}(\lambda,\tau,i)$ do not depend on $i$ and we can get them by dividing (23) by $\dim\tau$.
Let us evaluate the ingredients of (23). $(S_{N},S_{N-\eta}\times S_{\eta})$ is a Gelfand pair, which means that each irreducible representation of $S_{N}$ has at most one $S_{N-\eta}\times S_{\eta}$–invariant vector. Interpreting summations over $\tilde{g}\in S_{N-\eta}\times S_{\eta}$ in (21) as projections on invariants, we identify $\chi^{\tau}[\mathcal{A};\mathbf{n}]$ with spherical functions corresponding to this invariant vector. The latter spherical function is evaluated in [Dun78, Corollary 2.2], [Sta84, (3.15)], [CSST08, Theorems 6.1.10, 6.2.3] in terms of hypergeometric function ${}_{3}F_{2}$, which can also be identified with the Hahn polynomial. The value at length $2$ cycle is given as
@@
(1a), (2a), (3a), and (4a) are all instances of Theorem 3.1 or Theorem 5.4 with $\tau=(3)$. (2ab) and (3ab) are instances of Corollary 5.7. (4a) and (4c) are instances of Theorem 5.1 for $\tau=(3)$ and $\tau=(1,1,1)$, respectively. The sum of all four eigenvalues in (4b) matches Theorem 5.1 for $\tau=(2,1)$, where we have $\dim\tau=\chi^{\tau}(\mathrm{Id})=2$, $\chi^{\tau}\bigl{(}(1,2)\bigr{)}=0$, $\chi^{\tau}\bigl{(}(1,2,3)\bigr{)}=-1$. The individual eigenvalues in (4b) are not computed by any of our theorems; the formula (33) can be obtained by evaluating the matrices of the operators $T_{1}^{m}$, $T_{2}^{m}$, $T_{3}^{m}$ by diagonalization and then direct computation of the eigenvalues of $T_{1}^{m}+T_{2}^{m}+T_{3}^{m}$ (we can either deal with $6\times 6$ matrices in the space of all permutations of $x_{1}^{a}x_{2}^{b}x_{3}^{c}$ or $4\times 4$ matrices in the subspace spanned by (34)).
@@ 7. Appendix: Basic properties of $\mathcal{P}_{m}$
Operators $\mathcal{P}_{m}$, $m=1,2,\dots$, commute.
@@
In order to prove (35) for $a\neq b$, plug the definition of $\Delta_{a}$ into $\Delta_{a}\Delta_{b}$, open the parenthesis and note that the double sum splits into three types of summands: involving two, three, and four variables $x_{i}$. The parts involving two and four variables are obviously the same in $\Delta_{a}\Delta_{b}$ and in $\Delta_{b}\Delta_{a}$. The part involving three variables, $x_{a}$, $x_{b}$, $x_{j}$ is more tricky, it is:
@@
(37) $[x_{a},\Delta_{b}]=x_{a}\Delta_{b}-\Delta_{b}x_{a}=(a,b).$
@@
(39) $[x_{a}\mathcal{D}_{a},x_{b}\mathcal{D}_{b}]=\theta(x_{a}\mathcal{D}_{a}-x_{b}\mathcal{D}_{b})\cdot(a,b).$
@@
(40) $[x_{a}\mathcal{D}_{a},(x_{b}\mathcal{D}_{b})^{m}]=\sum_{k=0}^{m-1}(x_{b}\mathcal{D}_{b})^{k}[x_{a}\mathcal{D}_{a},x_{b}\mathcal{D}_{b}](x_{b}\mathcal{D}_{b})^{m-k-1}\\ =\theta\sum_{k=0}^{m-1}(x_{b}\mathcal{D}_{b})^{k}(x_{a}\mathcal{D}_{a}-x_{b}\mathcal{D}_{b})\cdot(a,b)(x_{b}\mathcal{D}_{b})^{m-k-1}=\\ \theta\sum_{k=0}^{m-1}(x_{b}\mathcal{D}_{b})^{k}(x_{a}\mathcal{D}_{a}-x_{b}\mathcal{D}_{b})\cdot(x_{a}\mathcal{D}_{a})^{m-k-1}\cdot(a,b)=\theta\bigl{(}(x_{a}\mathcal{D}_{a})^{m}-(x_{b}\mathcal{D}_{b})^{m}\bigr{)}\cdot(a,b).$
@@
$[(x_{a}\mathcal{D}_{a})^{l},(x_{b}\mathcal{D}_{b})^{m}]=\sum_{k=0}^{l-1}(x_{a}\mathcal{D}_{a})^{k}[x_{a}\mathcal{D}_{a},(x_{b}\mathcal{D}_{b})^{m}](x_{a}\mathcal{D}_{a})^{l-k-1}\\ =\theta\left(\sum_{k=m}^{m+l-1}(x_{a}\mathcal{D}_{a})^{m+k}(x_{b}\mathcal{D}_{b})^{m+l-k-1}-\sum_{k=0}^{l-1}(x_{a}\mathcal{D}_{a})^{k}(x_{b}\mathcal{D}_{b})^{m+l-k-1}\right)\cdot(a,b).$
Interchanging $l$ and $m$ and using the skew-symmetry $[f,g]=-[g,f]$, we also have
$[(x_{b}\mathcal{D}_{b})^{l},(x_{a}\mathcal{D}_{a})^{m}]=\theta\left(\sum_{k=0}^{m-1}(x_{a}\mathcal{D}_{a})^{k}(x_{b}\mathcal{D}_{b})^{m+l-k-1}-\sum_{k=l}^{m+l-1}(x_{a}\mathcal{D}_{a})^{l+k}(x_{b}\mathcal{D}_{b})^{m+l-k-1}\right)\cdot(a,b).$
@@
Changing the variables $x_{a}=\exp(\mathbf{i}z_{a})$, conjugating with $\prod_{a<b}\sin^{\theta}\left(\frac{z_{a}-z_{b}}{2}\right)$ and shifting by a constant, (42) turns into $-H$, where $H$ is the trigonometric version of the so-called Calogero-Moser-Sutherland Hamiltonian:
@@
Since $01-(i,j)=$ 0 on symmetric functions, $(x_{i}\mathcal{D}_{i})^{2}$ restricts as
@@
The permutation $(i,j)$ acts identically on symmetric functions and can be removed from the right-hand side of the last formula. Summing over all $i$, we arrive at (42). ∎
@@
As in the proof of Theorem 3.1, we note that the linear space $\bigcup\limits_{\mu\mid\mu\preceq\lambda}V_{\mu}$ is invariant both for $\mathcal{P}_{m}$ and for the symmetric group $S_{N}$. Because $S_{N}$–action commutes with $\mathcal{P}_{m}$ action, $\mathcal{P}_{m}$ preserves the subspace of symmetric polynomial inside $\bigcup\limits_{\mu\mid\mu\preceq\lambda}V_{\mu}$. Hence, $\mathcal{P}_{m}$ must have some symmetric eigenfunctions in $\bigcup\limits_{\mu\mid\mu\preceq\lambda}V_{\mu}$ and then by triangularity consideration, the unique symmetric eigenfunction with leading monomial in $V_{\lambda}$ has to be $F_{\lambda,(N),1}$. We conclude that $F_{\lambda,(N),1}$ is symmetric. Since $\mathcal{P}_{m}$ is self-adjoint by Lemma 4.7, its eigenfunctions are orthogonal with respect to scalar product of that lemma. We conclude that as $\lambda$’s vary, the functions $F_{\lambda,(N),1}$ form an orthogonal basis of the space $\Lambda_{N}$ of symmetric functions in $N$ variables and this basis is related to monomial symmetric functions $m_{\lambda}$ by a triangular transformation. Being such an orthogonal basis is a property which uniquely determines Jack polynomials and can be taken as one of their equivalent definitions, see [Sta89] or [Mac95, Section 6.10] and note that they use $\alpha=1/\theta$ as the parameter. ∎

### v2

@@ 5.4. $a\dots ab\dots b$ case with $p=2$
Another situation where the general formula of Theorem 5.4 is fully explicit is $p=2$. In fact, in this special case we are able to compute all eigenvalues $\mathrm{eig}_{m}(\lambda,\tau,i)$ rather than only their sums over $i$. We fix $N=1,2,\dots$ and $1\leq\eta\leq N-1$ and set the multiplicities to be $(n_{1},n_{2})=(N-\eta,\eta)$. We further choose $a>b\geq 0$ and set
@@
We need some facts from the representation theory of symmetric groups, which can be found e.g. in [Dun78], or [Sta84, Section 3], or [CSST08, Sections 6.1, 6.2]. In particular, $V_{\lambda}$ expands into a direct multiplicity one sum of irreducible representations $\tau=(N-k,k)$, $0\leq k\leq\min(\eta,N-\eta)$. The operator $(T_{1}^{m}+\dots+T_{N}^{M})$ commutes with the action of the symmetric group. Therefore, by Schur’s lemma, it necessarily acts as a multiple of the identity in each $V_{\lambda;\tau}$ (which would not have been true if $\tau$ had multiplicity larger than $1$ in $\lambda$). Hence, the eigenvalues $\mathrm{eig}_{m}(\lambda,\tau,i)$ do not depend on $i$ and we can get them by dividing (23) by $\dim\tau$.
Let us evaluate the ingredients of (23). $(S_{N},S_{N-\eta}\times S_{\eta})$ is a Gelfand pair, which means that each irreducible representation of $S_{N}$ has at most one $S_{N-\eta}\times S_{\eta}$-invariant vector. Interpreting summations over $\tilde{g}\in S_{N-\eta}\times S_{\eta}$ in (21) as projections on invariants, we identify $\chi^{\tau}[\mathcal{A};\mathbf{n}]$ with spherical functions corresponding to this invariant vector. The latter spherical function is evaluated in [Dun78, Corollary 2.2], [Sta84, (3.15)], [CSST08, Theorems 6.1.10, 6.2.3] in terms of the hypergeometric function $_{3}F_{2}$, which can also be identified with the Hahn polynomial. The value at length $2$ cycle is given as
@@
5.5. More complicated isotypes
In the follow-up work, the spherical functions $\chi\left[\mathcal{A},\mathbf{n}\right]$ have been computed for the hook isotypes $\left[N-b,1^{b}\right]$ for $p>b$ in [Dun25a] and for the isotypes $\left[N-k,k\right]$ for $p=3$ in [Dun25b].
@@ 5.5. More complicated isotypes
(i.a), (ii.a), (iii.a), and (iv.a) are all instances of Theorem 3.1 or Theorem 5.4 with $\tau=(3)$. (ii.ab) and (iii.ab) are instances of Corollary 5.7. (iv.a) and (iv.c) are instances of Theorem 5.1 for $\tau=(3)$ and $\tau=(1,1,1)$, respectively. The sum of all four eigenvalues in (iv.b) matches Theorem 5.1 for $\tau=(2,1)$, where we have $\dim\tau=\chi^{\tau}(\mathrm{Id})=2$, $\chi^{\tau}\bigl{(}(1,2)\bigr{)}=0$, $\chi^{\tau}\bigl{(}(1,2,3)\bigr{)}=-1$. The individual eigenvalues in (iv.b) are not computed by any of our theorems; the formula (32) can be obtained by evaluating the matrices of the operators $T_{1}^{m}$, $T_{2}^{m}$, $T_{3}^{m}$ by diagonalization and then direct computation of the eigenvalues of $T_{1}^{m}+T_{2}^{m}+T_{3}^{m}$ (we can either use $6\times 6$ matrices in the space of all permutations of $x_{1}^{a}x_{2}^{b}x_{3}^{c}$ or $4\times 4$ matrices in the subspace spanned by (33)).
@@ 7. Appendix: Basic properties of $\mathcal{P}_{m}$
The operators $\mathcal{P}_{m}$, $m=1,2,\dots$, mutually commute.
@@
In order to prove (34) for $a\neq b$, we plug the definition of $\Delta_{a}$ into $\Delta_{a}\Delta_{b}$ and get
$\sum_{j\neq a}\sum_{j^{\prime}\neq b}\frac{1-(a,j)}{x_{a}-x_{j}}\cdot\frac{1-(b,{j^{\prime}})}{x_{b}-x_{j^{\prime}}}.$
Note that the double sum splits into three types of summands: involving two, three, and four distinct variables $x_{i}$. The parts involving two and four variables are obviously the same in $\Delta_{a}\Delta_{b}$ and in $\Delta_{b}\Delta_{a}$. The part involving three variables, $x_{a}$, $x_{b}$, $x_{j}$ is more tricky, it is:
@@
(36) $x_{a},\Delta_{b}]=x_{a}\Delta_{b}-\Delta_{b}x_{a}=(a,b).$
@@
(38) $x_{a}\mathcal{D}_{a},x_{b}\mathcal{D}_{b}]=\theta(x_{a}\mathcal{D}_{a}-x_{b}\mathcal{D}_{b})\cdot(a,b).$
@@
(39) $x_{a}\mathcal{D}_{a},(x_{b}\mathcal{D}_{b})^{m}]=\sum_{k=0}^{m-1}(x_{b}\mathcal{D}_{b})^{k}[x_{a}\mathcal{D}_{a},x_{b}\mathcal{D}_{b}](x_{b}\mathcal{D}_{b})^{m-k-1}\\ =\theta\sum_{k=0}^{m-1}(x_{b}\mathcal{D}_{b})^{k}(x_{a}\mathcal{D}_{a}-x_{b}\mathcal{D}_{b})\cdot(a,b)(x_{b}\mathcal{D}_{b})^{m-k-1}=\\ \theta\sum_{k=0}^{m-1}(x_{b}\mathcal{D}_{b})^{k}(x_{a}\mathcal{D}_{a}-x_{b}\mathcal{D}_{b})\cdot(x_{a}\mathcal{D}_{a})^{m-k-1}\cdot(a,b)=\theta\bigl{(}(x_{a}\mathcal{D}_{a})^{m}-(x_{b}\mathcal{D}_{b})^{m}\bigr{)}\cdot(a,b).$
@@
$(x_{a}\mathcal{D}_{a})^{l},(x_{b}\mathcal{D}_{b})^{m}]=\sum_{k=0}^{l-1}(x_{a}\mathcal{D}_{a})^{k}[x_{a}\mathcal{D}_{a},(x_{b}\mathcal{D}_{b})^{m}](x_{a}\mathcal{D}_{a})^{l-k-1}\\ =\theta\left(\sum_{k=m}^{m+l-1}(x_{a}\mathcal{D}_{a})^{m+k}(x_{b}\mathcal{D}_{b})^{m+l-k-1}-\sum_{k=0}^{l-1}(x_{a}\mathcal{D}_{a})^{k}(x_{b}\mathcal{D}_{b})^{m+l-k-1}\right)\cdot(a,b).$
Interchanging $l$ and $m$ and using the skew-symmetry [f,g]=[g,f][f,g]=-[g,f], we also have
$(x_{b}\mathcal{D}_{b})^{l},(x_{a}\mathcal{D}_{a})^{m}]=\theta\left(\sum_{k=0}^{m-1}(x_{a}\mathcal{D}_{a})^{k}(x_{b}\mathcal{D}_{b})^{m+l-k-1}-\sum_{k=l}^{m+l-1}(x_{a}\mathcal{D}_{a})^{l+k}(x_{b}\mathcal{D}_{b})^{m+l-k-1}\right)\cdot(a,b).$
@@
Changing the variables $x_{a}=\exp(\mathbf{i}z_{a})$, conjugating with $\prod_{a<b}\sin^{\theta}\left(\frac{z_{a}-z_{b}}{2}\right)$ and shifting by a constant, (41) turns into H-H, where $H$ is the trigonometric version of the so-called Calogero-Moser-Sutherland Hamiltonian:
@@
Since 1(i,j)=01-(i,j)=0 on symmetric polynomials, $(x_{i}\mathcal{D}_{i})^{2}$ restricts as
@@
The permutation $(i,j)$ acts identically on symmetric polynomials and can be removed from the right-hand side of the last formula. Summing over all $i$, we arrive at (41). ∎
@@
As in the proof of Theorem 4.3, we note that the linear space $\bigoplus\limits_{\mu\mid\mu\preceq\lambda}V_{\mu}$ is invariant both for $\mathcal{P}_{m}$ and for the symmetric group $S_{N}$. Because $S_{N}$-action commutes with $\mathcal{P}_{m}$ action, $\mathcal{P}_{m}$ preserves the subspace of symmetric polynomial inside $\bigoplus\limits_{\mu\mid\mu\preceq\lambda}V_{\mu}$. Hence, $\mathcal{P}_{m}$ must have some symmetric eigenfunctions in $\bigoplus\limits_{\mu\mid\mu\preceq\lambda}V_{\mu}$ and then by triangularity consideration, the unique symmetric eigenfunction with leading monomial in $V_{\lambda}$ has to be $F_{\lambda,(N),1}$. We conclude that $F_{\lambda,(N),1}$ is symmetric. Since $\mathcal{P}_{m}$ is self-adjoint by Lemma 4.7, its eigenfunctions are orthogonal with respect to scalar product of that lemma. We conclude that as $\lambda$’s vary, the polynomials $F_{\lambda,(N),1}$ form an orthogonal basis of the space $\Lambda_{N}$ of symmetric polynomials in $N$ variables and this basis is related to monomial symmetric polynomials $m_{\lambda}$ by a triangular transformation. Being such an orthogonal basis is a property which uniquely determines Jack polynomials and can be taken as one of their equivalent definitions, see [Sta89] or [Mac95, Section 6.10] and note that they use $\alpha=1/\theta$ as the parameter. ∎